OpenStudy (anonymous):

the sum of four times a number and 3 is equal to three times the number

OpenStudy (anonymous):

let the number be x 4times the number=4x 3times the number= 3x therefore 4x+3=3x collect like terms 4x-3x=-3 x=-3 the number is -3

OpenStudy (anonymous):

five times a number subtracted from twelve is triple is the number

OpenStudy (anonymous):

12-5x=3x collect like terms 12=5x+3x 12=8x 12/8=x so the number is 3/2 or 1.5

OpenStudy (anonymous):

if the sum of a number and one is tripled the result is five less than twice the number